Coulomb's Law and Net Electric Field

Solve the following word problems. Show complete solutions to your answer.

1. Find the electronic force between charges of +2.0 C and +5.0 C separated by distance of 75 m in a vacuum.

2. Two charge of +8.0 mC and -6.0 mC attract each other with a force of 3.0 x 103 N in a vacuum. What is the distance between the charges?

3. A charge of 12 uC when placed in electric field experiences a force of 648 N. What is the magnitude of the electric field strength?

4. What happens to the force between two charges metal spheres in a vacuum if the charge on each is doubled and the distance between them is multiplied by three?


1
Expert's answer
2021-03-26T11:37:18-0400

1)


"F=9\cdot10^9\frac{2\cdot 5}{75^2}=1.6\cdot10^7\ N"

2)


"9\cdot10^9\frac{0.008\cdot 0.006}{d^2}=3\cdot10^3\\d=12\ m"

3)


"E=\frac{648}{12\cdot10^{-6}}=5.4\cdot10^{7}\frac{N}{C}"

4)


"F=k\frac{q^2}{d^2}\\F'=k\frac{4q^2}{9d^2}=\frac{4}{9}F"